OpenStudy (anonymous):

Integrate log { x-6/x-1 Plz help

geerky42 (geerky42):

Like this? \[\int \log\left(\dfrac{x-6}{x-2}\right)dx\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Sorry i am on my ipad so cant write proper words

OpenStudy (anonymous):

The answer at the back is x-5 ln (x-1) +c

OpenStudy (anonymous):

I dunno how to get that

geerky42 (geerky42):

can you check again? I don't think answer is x-5 ln (x-1) + C

OpenStudy (anonymous):

The question is Show that x-6/x-1= 1- 5/x-1 Theh it says hence find { x-6/x-1

OpenStudy (anonymous):

And answer at the back is x-5 loge (x-1) +c

geerky42 (geerky42):

ah I see, for first part, you have to use polynomial long division.

OpenStudy (aravindg):

Or simply split the integral using propertt log a/b=log a-log b

OpenStudy (aravindg):

*propery

OpenStudy (anonymous):

But how thats gonna help

OpenStudy (aravindg):

Because integral of log (x-a)=x(log(x-a)-1)

OpenStudy (anonymous):

Do u mind solving it

OpenStudy (aravindg):

I don't think I should. I just gave you the formula. Plug in and get your answer.

OpenStudy (anonymous):

Honestly i dunno hw that formula apply to fraction

OpenStudy (aravindg):

First use log a/b=log a-log b Then apply formula for each term.

OpenStudy (anonymous):

So ∫ x/6 / x/1

OpenStudy (anonymous):

i think your question is \[show ~that ~\frac{ x-6 }{ x-1 }=1-\frac{ 5 }{ x-1 } and~hence~integrate~\it.\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{ x-6 }{ x-1 }=\frac{ x-1-5 }{ x-1 }=\frac{ x-1 }{ x-1 }-\frac{ 5 }{ x-1 }=?\]

OpenStudy (anonymous):

I know how to prove the equation its the integrate answer

OpenStudy (anonymous):

I dont get where x-5 comes from in x-5 ln (x-1 )+ answer

OpenStudy (anonymous):

\[\frac{ x-1 }{ x-1 }-\frac{ 5 }{ x-1 }=1-\frac{ 5 }{ x-1 }\] \[\int\limits \frac{ x-6 }{ x-1 }dx=\int\limits \left( 1-\frac{ 5 }{ x-1 } \right)dx=\int\limits dx-5 \int\limits \frac{ dx }{ x-1 }+c\] =?